OpenStudy (anonymous):

I'm trying to figure out the asymptotes for y= -2 sec (pi*x/3) ? with the period being 6. By looking at the graph it kind of looks like 1.5 and 4.5 but I'm confused as to how to find the exact x value?

OpenStudy (anonymous):

hey sourwing hows it goin.. do you think you may be able to point me in the right direction here?

OpenStudy (anonymous):

what do you think sec(u) is?

OpenStudy (anonymous):

I just set pi*x/3 equal to pi/2 and 3pi/2, since the cos of theta is at pi/2 and 3pi/2 when it's zero, but I'm doing something wrong?

OpenStudy (anonymous):

@sourwing is wierd

OpenStudy (anonymous):

x)

OpenStudy (anonymous):

lol @Pooch @lahernan why did you set xpi/3 = pi/2 and 3pi/2?

OpenStudy (anonymous):

I know sec is = 1/cos theta

OpenStudy (anonymous):

good, what did you do next?

OpenStudy (anonymous):

well, since the function of sec is just the reciprocal of cos I just set pi*x/3=pi/2 to figure the the asymptote is when cos can't be zero, and cos is zero on the unit circle at pi/2 ? but I'm lost somewhere because i know for whatever reason that's not right?

OpenStudy (anonymous):

You're on the right track. Except you forgot the sec(x) is a periodic function. Which means has how many asymoptes?

OpenStudy (anonymous):

oh yeah. sorry. Infinite. I guess I should of said, just for one period. Period being equal to 6

OpenStudy (anonymous):

Do you you have the graph. There are infinitely many parts of the graph that has the period of 6.

OpenStudy (anonymous):

umm , i have it on my computer, but i dont know how to post it on here?

OpenStudy (anonymous):

What is the domain on the graph?

OpenStudy (anonymous):

So how are you supposed to know which period to designate?

OpenStudy (anonymous):

domain is all real numbers except for according to what I'm thinking is pi/2+pi?

OpenStudy (anonymous):

well yes, But I was preferring the domain on the graph. Is it like x = 0 to x = 6 ?

OpenStudy (anonymous):

oh yes. correct.

OpenStudy (anonymous):

well, x = 3/2 and x = 9/2 are the vertical asymptotes on that specific range. Unless the questions wants you to find ALL vertical asymptotes

OpenStudy (anonymous):

no those are the only ones. but I don't understand how did you get those values?

OpenStudy (anonymous):

well if 6 is the period, what do you think the value of the function at x= 1/2 + 6 is? or at 9/2 + 6?

OpenStudy (anonymous):

it would be the same value just adding 6.

OpenStudy (anonymous):

typo at x = 3/2 + 6

OpenStudy (anonymous):

Yes, it would be the same, which means f is also what at 3/2 + 6?

OpenStudy (anonymous):

zero

OpenStudy (anonymous):

yes, 0 for cos(x) but what about 1/cos(x)?

OpenStudy (anonymous):

but i'm trying to think in terms algebraically? as if i had no graph

OpenStudy (anonymous):

Yes, i'm trying to help you get all asymtopes. So what is the value of the original function at x = 3/2 + 6 or at x = 9/2 + 6?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

no, we're looking for *vertical asymptotes* yes? which means the function is what at those asymptotes?

OpenStudy (anonymous):

oh you mean in radians?

OpenStudy (anonymous):

the function is unbounded?

OpenStudy (anonymous):

well, x is already in radian. The answer I was looking for is the function is *undefined* at those vertical asymptotes

OpenStudy (anonymous):

see thats exactly where I'm mixed up somehow because if the value for 1/cos is undefined at cos =0 why don't I get those two values 3/2 and 9/2 when I set pix/3= pi/2?

OpenStudy (anonymous):

well, let me give you a simple function. Say f(x) = 1/(2x). Clearly the function is undefined at x = 0 yes?

OpenStudy (anonymous):

sorry bout that. yes

OpenStudy (anonymous):

which means x ≠ 0 yes?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Back the the original problem. -2sec(pix/3) means -2/cos(pix/3) which means cos(pix/3) ≠ 0, yes?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

ok. so it would be negative two divided by cos pi/3 which would be 1/2?

OpenStudy (anonymous):

i have to constanly refresh my browser i keep losing connection.

OpenStudy (anonymous):

no, -2 is on the numerator so there is no problem. when you were solving cos(pix/3) ≠ 0, you were really trying to solve cos(pix/3) = 0. Whatever the solutions are, those are the value x can *not* be yes? so really x ≠ 3/2 and x ≠ 9/2. But the definition of vertical asymptote is where x can *not* be. Which means vertical asymptotes are at x = 3/2, x = 9/2 yes?

OpenStudy (anonymous):

which also means the function is *undefined* at those x values yes? Now the only problem we have left is how to find the other ones since there are infinitely many of them

OpenStudy (anonymous):

ok. i see what your saying. So I was trying to solve for the problem when there was none.

OpenStudy (anonymous):

man, i think im so used to dealing with pi, i'm not connecting the dots, when i see numbers. thank you for your help . As always, you've shown me the path. Thanks.

OpenStudy (anonymous):

so when i asked you what is the valued at 3/2 + 6. You should have said *undefined*, yes?

OpenStudy (anonymous):

"":) yea

OpenStudy (anonymous):

what do you think the value of the function at x = 3/2 + 2(6) is?

OpenStudy (anonymous):

undefined

OpenStudy (anonymous):

good, how about at x = 3/2 + 3(6) ?

OpenStudy (anonymous):

undefined. Yea i need a personal tutor, just to really get the basics to sink in. Thanks Sour

OpenStudy (anonymous):

no problem. What do you the final answer is?

OpenStudy (anonymous):

in other works, where do you think all the asymptotes are at?

OpenStudy (anonymous):

*words*, ..

OpenStudy (anonymous):

where the function is undefined. thanks man. I'll definitely hit u up next time i get on. TTyl.